# High-dimensional dynamics in low-dimensional networks

### v1

@@ Acknowledgments
is the sampled value of $R$ (assuming $J^{{\bf{x}}}=I$) and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (24) gives a much more accurate prediction (Supplementary Figure S.13a, green dots) This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (23) was accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
As concluded at the end of the previous section, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
Combining these conclusions with those reached in Section S.2.3, we can summarize as follows
Low-rank suppression (as in Figure 1c) occurs whenever $Q=I-W$ has at least one asymptotically large singular value.
High-dimensional dynamics (as in Figure 1d) occur whenever $Q=I-W$ lacks any asymptotically small singular values.
In the previous section, we showed that $Q=I-W$ has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank. Conditions under which $Q=I-W$ lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
@@
which measures the alignment between the left and right singular vectors, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$. Specifically, $P_{jk}={\bf{v}}_{k}\cdot{\bf{u}}_{k}$ so that $P_{jk}=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$. Note that singular values of $P$ are bounded by unity, $\sigma_{P}\leq 1$. We next show that if all singular values of $P$ are $\mathcal{O}(1)$ then the network exhibits high-dimensional responses to high-dimensional inputs.
@@
Under the model and assumptions of Sections S.2.2 and S.2.4, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
@@
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever $Q=I-W$ has an asymptotically small singular value. Therefore, our original claim is equivalent to the following:
If $Q=I-W$ has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
@@
The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2g. Since ${\bf{z}}$ is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$. Indeed, in Figure 2g, the angle between ${\bf{u}}_{2}$ and the dominant principle component direction of ${\bf{z}}$ is $18^{\circ}$.
@@
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (20) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (22). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (20) instead of the identity itself, as in Eq. (22). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$).

### v2

@@ Acknowledgments
is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (17) gives a much more accurate prediction (Supplementary Figure S.13a, green dots). This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (16) was more accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics can only occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
Combining these conclusions with those in Claim 1, we can summarize as follows
Low-rank suppression (as in Figure 1g) occurs whenever Q=IWQ=I-W has at least one asymptotically large singular value.
High-dimensional dynamics (as in Figure 1d) can only occur whenever Q=IWQ=I-W lacks any asymptotically small singular values.
Previously, in proving Claim 1, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank. Conditions under which Q=IWQ=I-W lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
@@
which measures the alignment between the left and right singular vectors, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$. Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ so that $|P_{jk}|=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$. Note that singular values of $P$ are bounded by unity, $\sigma_{P}\leq 1$. We next show that if all singular values of $P$ are $\mathcal{O}(1)$ then the network exhibits high-dimensional responses to high-dimensional inputs.
@@
Under the model and assumptions 1–3 in the main text, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
@@
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2 and Figure 3g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever Q=IWQ=I-W has an asymptotically small singular value. Therefore, our original claim is equivalent to the following:
If Q=IWQ=I-W has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
@@
The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2 and Figure 3g. Since $\bf{z}$} is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$.
@@
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (13) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (15). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (13) instead of the identity itself, as in Eq. (15). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$). Next, in Supplementary Text S.3, we generalize away from the assumption of slow inputs using a Laplace transform.
S.3 The analysis of fast and transient dynamics.
Above, we considered slow or static perturbations, ${\bf{x}}(t)$. Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models). This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics. We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales. We can no longer rely on the quasi-steady state approximation from Eq. (2) in this case. Instead, we take the Laplace transform in Eq. (6) to obtain
$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (18)
where $\hat{\psi}(s)=\int\psi(t)e^{-st}dt$ is the Laplace transform of $\psi(t)$ and
$\hat{H}(s)$ $=(1+\tau s)^{-1}W$ (19)
$\hat{B}(s)$ $=(1+\tau s)^{-1}.$
Solving the implicit Eq. (18) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
$\hat{{\bf{z}}}=[I-\hat{H}]^{-1}\hat{B}\hat{{\bf{x}}}$. (20)
The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$ because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$. At mode $s=0$ (the static or “DC” mode), we recover the connectivity from the steady-state or quasi-steady state analysis, $\hat{H}(0)=W$. Similarly, $\hat{B}(0)=1$. Hence, the quasi-steady state analysis from above evaluated how the DC component of the input, $\hat{{\bf{x}}}(0)$, affects the DC component of the dynamics, $\hat{{\bf{z}}}(0)$, even when we do not assume that ${\bf{x}}(t)$ changes slowly. The analysis here is a generalization of the quasi-steady state analysis to non-static modes, $s\neq 0$.
Note also that Eq. (13) from Supplementary Text S.2 can be derived from Eqs. (20) and (19) above by first taking $s=-2\pi if$ in Eq. (20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the approach in this section is a generalization of the approach in Supplementary Text S.2.
To interpret Eq. (20), let us consider two separate limits. At modes that are much slower than the intrinsic timescale of the dynamics, $|s|\ll 1/\tau$, we have that $\hat{H}(s)\approx W$ and $\hat{B}(s)\approx 1$, so
$\hat{{\bf{z}}}(s)\approx[I-W]^{-1}\hat{{\bf{x}}}(s)$
and we recover the results from our quasi-steady state analysis at these modes. In particular, Claim 1 holds and we expect low-rank suppression at mode $s$ whenever $W$ is strongly low rank. While this may at first seem like a repetition of our results from the quasi-steady state analysis, note that we do not need to assume that ${\bf{x}}(t)$ varies slowly here. The Laplace analysis at slow modes ($|s|\ll 1/\tau$) applies even when ${\bf{x}}(t)$ has power at slow and fast modes.
At fast modes, $|s|\gg\tau$, the identity term in Eq. (20) dominates the $\hat{H}$ and $\hat{B}$ terms so that
$\hat{{\bf{z}}}(s)\approx\hat{{\bf{x}}}(s)$
at fast modes. Therefore, fast modes (representing transient responses to transient input dynamics) do not exhibit low-rank suppression, but they do produce high-dimensional responses to high-dimensional stimuli. Modes between these two extremes interpolate between them.
In summary, when our previous assumption that ${\bf{x}}(t)$ varies more slowly than $\tau$ is violated, we expect to see low-rank suppression at slow temporal modes, but not at fast temporal modes.
Additionally, working in the Laplace domain allows us to expand our class of models. In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq. (4), which we later linearized. The dynamics in Eq. (4) capture a large class of models, but are not fully general. Specifically, in Eq. (4), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$. However, some systems like integral equations, integro-differential equations, operator equations, and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq. (4). In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (21)
In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivariant, causal function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$. Eq. (21) encompasses an extremely broad class of systems, including those described by Eq. (4).
A first order Volterra expansion gives Eq. (18) as an implicit equation for ${\bf{z}}(t)$, but the matrices $\hat{H}$ and $\hat{B}$ are no longer defined as above. Instead, the matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline. Under the dynamics in Eq. (6), we recover the $\hat{H}$ and $\hat{B}$ defined in Eq. (19). Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq. (21).